If X has uniform distribution over [0,1], find the density function and distribution function of Y D 3X: Identify the distribution of Y


1
Expert's answer
2021-06-11T16:34:21-0400

Solution:

"X\sim U[0,1]"

"Y=3X
\\\Rightarrow X=\dfrac Y3"

"P(Y\le r)=P[0\le X\le \dfrac r3]=\dfrac r3" ...(i)

As "0<x<1, 0<\dfrac Y3<1"

"\Rightarrow 0<Y<3"

Now, differentiating (i) to get its PDF.

"f_Y(t)=(\dfrac t3)'=\dfrac 13"

Thus, distribution function:

"P(Y\le r)=\{\dfrac r3, 0<r<3
\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0,\ \ \ otherwise"

And density function:

"f_Y(t)=\dfrac 13;0<t<3
\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0,\ \ \ otherwise"